How to solve:
-3/4x - 1/2 = 5/8

Answer by algebrahouse.com(1659)

(-3/4)x - 1/2 = 5/8
(-3/4)x = 9/8 {added 1/2 to both sides by getting common denominator}
x = 9/8(-4/3) {divided both sides by -3/4, multiplying by reciprocal, -4/3}
x = -3/2 {cancelled and multiplied what's left}

For more help from me, visit: www.algebrahouse.com